Evaluate the following limit:
limt→1−t2−|t−1|−1|t−1|.
1
Expert's answer
2020-05-26T20:01:23-0400
"\lim\limits_{t\to1}(-{t^2-|t-1|-1\over |t-1|})"

"\lim\limits_{t\to1^+}(-{t^2-|t-1|-1\over |t-1|})=\lim\limits_{t\to1^+}(-{t^2-t+1-1\over t-1})="

"=\lim\limits_{t\to1^+}(-{t(t-1)\over t-1})=-\lim\limits_{t\to1^+}t=-1"


"\lim\limits_{t\to1^-}(-{t^2-|t-1|-1\over |t-1|})=\lim\limits_{t\to1^-}({t^2+t-1-1\over t-1})="

"=\lim\limits_{t\to1^-}{(t+2)(t-1)\over t-1}=\lim\limits_{t\to1^-}(t+2)=3"

"\lim\limits_{t\to1^+}(-{t^2-|t-1|-1\over |t-1|})=-1\not=3=\lim\limits_{t\to1^-}(-{t^2-|t-1|-1\over |t-1|})"

Therefore


"\lim\limits_{t\to1}(-{t^2-|t-1|-1\over |t-1|})"

does not exist.